# A mesh-free hybrid Chebyshev-Tucker tensor format with applications to multi-particle modelling

### v1

@@ 2.2 Transforming hybrid representation to the standard low-rank Tucker format
Let the two-level third order orthogonal Tucker tensor ${\bf V}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ be given in the form Eq. 2.3, where the core coefficient is parametrized by the rank ${\bf r}^{0}=(r_{1}^{0},r_{2}^{0},r_{3}^{0})$ orthogonal Tucker tensor
$\bm{\beta}=\bm{\beta}^{0}\times_{1}G^{(1)}\times_{2}G^{(2)}\times_{3}G^{(3)}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}},$ (2.9)
with the core tensor $\bm{\beta}^{0}\in\mathbb{R}^{r_{1}^{0}\times r_{2}^{0}\times r_{3}^{0}}$. Then the initial tensor ${\bf V}$ takes the form of a rank ${\bf r}^{0}=(r_{1}^{0},r_{2}^{0},r_{3}^{0})$ orthogonal Tucker representation
${\bf V}=\bm{\beta}^{0}\times_{1}(V^{(1)}G^{(1)})\times_{2}(V^{(2)}G^{(2)})\times_{3}(V^{(3)}G^{(3)}),$ (2.10)
with the core tensor $\bm{\beta}^{0}$ and the orthogonal Tucker factor matrices $V^{(\ell)}G^{(\ell)}\in\mathbb{R}^{n_{\ell}\times r_{\ell}^{0}}$.
Proof.
Substituting Eq. 2.9 into Eq. 2.3 and applying contracted product operations to the resultant Tucker tensor
${\bf V}=[\bm{\beta}^{0}\times_{1}G^{(1)}\times_{2}G^{(2)}\times_{3}G^{(3)}]\times_{1}V^{(1)}\times_{2}V^{(2)}\times_{3}V^{(3)},$ (2.11)
proves the orthogonality $V^{(\ell)}G^{(\ell)}[G^{(\ell)}]^{T}[V^{(\ell)}]^{T}=I_{n_{\ell}}$, which completes the proof. ∎
Notice that if Eq. 2.9 represents a non-orthogonal Tucker tensor, then $\{V^{(\ell)}G^{(\ell)}\}$ appear to be the non-orthogonal $n_{\ell}\times r_{\ell}^{0}$ matrices, also that the resultant conventional Tucker tensor (2.10) can be easily transformed to the canonical tensor format by using the so-called Tuck-to-can transform, see [31, Def. 2.3, (2.6)], [28, Rem. 2.7], and [26]. This Tuck-to-can algorithm will be applied to the core tensor $\bm{\beta}^{0}$ in Eq. 2.9.
The following statement is important in further use of the hybrid functional representation in multilinear algebra, see also (3.9).
@@
The representation Eq. 2.7 can be converted to the standard functional Tucker format as in Definition 2.1, i.e., Eq. 2.7 can be equivalently written as
$f(x_{1},x_{2},x_{3})=\bm{\beta}\times_{1}v^{(1)}(x_{1})\times_{2}v^{(2)}(x_{2})\times_{3}v^{(3)}(x_{3}),$ (2.12)
where the Tucker factor functions are given by $v^{(\ell)}(x_{\ell})=T_{0:m_{\ell}-1}(x_{\ell})V^{(\ell)}\in\mathbb{R}^{r_{\ell}}$, i.e., each entry of $v^{(\ell)}$ is a linear combination of the Chebyshev polynomials. If Eq. 2.9 present the orthogonal Tucker tensor, then $\{T_{0:m_{\ell}-1}(x_{\ell})V^{(\ell)}\}$ appear to be the set of $r_{\ell}^{0}$ orthogonal polynomials.
The proof of this proposition is in essence the same as the proof of Proposition 2.4, i.e., follows directly from the definition of contracted product operations $\times_{\ell}$. However, here we provide a detailed proof to write out several useful equivalent representations of the ChebTuck format explicitly.
Proof.
From Eq. 2.7 we have
$f(x_{1},x_{2},x_{3})$ $={\bf C}\times_{1}T_{0:m_{1}-1}(x_{1})\times_{2}T_{0:m_{2}-1}(x_{2})\times_{3}T_{0:m_{3}-1}(x_{3})$ (2.13)
$=\sum_{i_{1}=1}^{m_{1}}\sum_{i_{2}=1}^{m_{2}}\sum_{i_{3}=1}^{m_{3}}{\bf C}_{i_{1},i_{2},i_{3}}T_{i_{1}-1}(x_{1})T_{i_{2}-1}(x_{2})T_{i_{3}-1}(x_{3})$
$=\sum_{i_{1}=1}^{m_{1}}\sum_{i_{2}=1}^{m_{2}}\sum_{i_{3}=1}^{m_{3}}\sum_{j_{1}=1}^{r_{1}}\sum_{j_{2}=1}^{r_{2}}\sum_{j_{3}=1}^{r_{3}}\beta_{j_{1},j_{2},j_{3}}V^{(1)}_{i_{1},j_{1}}V^{(2)}_{i_{2},j_{2}}V^{(3)}_{i_{3},j_{3}}T_{i_{1}-1}(x_{1})T_{i_{2}-1}(x_{2})T_{i_{3}-1}(x_{3})$
$=\sum_{j_{1}=1}^{r_{1}}\sum_{j_{2}=1}^{r_{2}}\sum_{j_{3}=1}^{r_{3}}\beta_{j_{1},j_{2},j_{3}}\sum_{i_{1}=1}^{m_{1}}V^{(1)}_{i_{1},j_{1}}T_{i_{1}-1}(x_{1})\sum_{i_{2}=1}^{m_{2}}V^{(2)}_{i_{2},j_{2}}T_{i_{2}-1}(x_{2})\sum_{i_{3}=1}^{m_{3}}V^{(3)}_{i_{3},j_{3}}T_{i_{3}-1}(x_{3})$
@@
where $v^{(\ell)}_{j_{\ell}}(x_{\ell})=\sum_{i_{\ell}=1}^{m_{\ell}}V^{(\ell)}_{i_{\ell},j_{\ell}}T_{i_{\ell}-1}(x_{\ell})$ for $\ell=1,2,3$. The desired result then follows. The orthogonality property is based on the argument analogues to the algebraic case but reformulated for polynomial arrays and note that the inner product should be properly weighted. ∎
@@
From the equivalent representations Eq. 2.13 of the ChebTuck format, we can see that it is a degree $(m_{1},m_{2},m_{3})$ polynomial in Chebyshev basis, where the coefficients are given as a Tucker tensor. For this reason, we sometimes also write a subscript ${\bf m}$ to denote the ChebTuck format of certain degrees, e.g., $f_{\bf m}$.
3 Numerical schemes for hybrid Chebyshev-Tucker approximation of trivariate functions
In what follows, we introduce the numerical schemes for rank-structured approximation of trivariate functions in the novel ChebTuck tensor format introduced in Definition 2.3. The basic idea is to first compute the Chebyshev coefficient tensor ${\bf C}$, either in full tensor format or in some low-rank format, and then convert ${\bf C}$ to the desired Tucker format with quasi-optimal $\varepsilon$-rank. Our techniques can be applied directly to multivariate functions with moderate regularity for which the Chebyshev tensor product interpolation demonstrates satisfactory convergence rates. Furthermore, we extend the presented approach beyond the class of regular functions and successfully apply it to functions with multiple singularities by using the range-separated tensor decomposition.
@@ 3.1.1 Functional case
Given a function $f:[-1,1]^{3}\to\mathbb{R}$, we consider its approximation by the tensor product of Chebyshev polynomials in the form (c.f. Eq. 2.7)
$f(x_{1},x_{2},x_{3})\approx\tilde{f}_{\bf m}(x_{1},x_{2},x_{3})=\sum_{i_{1}=1}^{m_{1}}\sum_{i_{2}=1}^{m_{2}}\sum_{i_{3}=1}^{m_{3}}{\bf C}_{i_{1},i_{2},i_{3}}T_{i_{1}-1}(x_{1})T_{i_{2}-1}(x_{2})T_{i_{3}-1}(x_{3}),$ (3.1)
where $T_{k}(x)$ is the $k$-th Chebyshev polynomial of the first kind, ${\bf C}\in\mathbb{R}^{m_{1}\times m_{2}\times m_{3}}$ is the Chebyshev coefficient tensor and ${\bf m}=(m_{1},m_{2},m_{3})$ is the degree of the Chebyshev polynomials in each mode.
There are mainly two ways to view the approximant $\tilde{f}_{\bf m}$ which leads to two different methods to compute the Chebyshev coefficient tensor ${\bf C}$. The first way is to view $\tilde{f}_{\bf m}$ as a truncation of the tensor product Chebyshev series expansion of $f$ and compute the coefficients by projection and integration. The analytic expressions for the Chebyshev coefficients in this case are straightforward to derive. In fact, each coefficient is a $w$-weighted inner product of $f$ and the corresponding Chebyshev basis, i.e.,
${\bf C}_{i_{1},i_{2},i_{3}}=\int_{[-1,1]^{3}}f(x_{1},x_{2},x_{3})T_{i_{1}-1}(x_{1})T_{i_{2}-1}(x_{2})T_{i_{3}-1}(x_{3})w(x_{1},x_{2},x_{3})\operatorname{d}x_{1}\operatorname{d}x_{2}\operatorname{d}x_{3},$ (3.2)
see e.g., [36, 37], but the computation is troublesome since one needs to evaluate those 3-dimensional integrals. The second way is to view $\tilde{f}_{\bf m}$ as an interpolation of $f$ at Chebyshev nodes, namely, $\tilde{f}_{\bf m}$ satisfies the following interpolation property
$\tilde{f}_{\bf m}(s_{i_{1}}^{(1)},s_{i_{2}}^{(2)},s_{i_{3}}^{(3)})=f(s_{i_{1}}^{(1)},s_{i_{2}}^{(2)},s_{i_{3}}^{(3)}),\quad\forall i_{\ell}=1,\cdots,m_{\ell},\quad\ell=1,2,3,$ (3.3)
where $s_{i_{\ell}}^{(\ell)}$ are the Chebyshev nodes of the second kind on each mode given by
$s_{i_{\ell}}^{(\ell)}=\cos\left(\frac{(i_{\ell}-1)\pi}{m_{\ell}-1}\right),\quad i=1,\cdots,m.$ (3.4)
The derivation of the Chebyshev coefficient tensor ${\bf C}$ in this case is more involved while it turns out that the computation can be done efficiently by the discrete cosine transform (DCT). The accuracy of these two approaches is normally similar [36] and we will therefore focus on the interpolation approach in this paper. See also [46, 16, 22].
@@
${\bf T}_{i_{1},i_{2},i_{3}}=f(s_{i_{1}}^{(1)},s_{i_{2}}^{(2)},s_{i_{3}}^{(3)}),\quad i_{\ell}=1,\cdots,m_{\ell},\quad\ell=1,2,3$ (3.5)
and the DCT matrix $W^{(\ell)}\in\mathbb{R}^{m_{\ell}\times m_{\ell}}$ as
$W^{(\ell)}=\dfrac{2}{m_{\ell}-1}\begin{pmatrix}\frac{1}{4}T_{0}(s_{1})&\frac{1}{2}T_{0}(s_{2})&\frac{1}{2}T_{0}(s_{3})&\dots&\frac{1}{4}T_{0}(s_{m_{\ell}})\\ \frac{1}{2}T_{1}(s_{1})&T_{1}(s_{2})&T_{1}(s_{3})&\dots&\frac{1}{2}T_{1}(s_{m_{\ell}})\\ \frac{1}{2}T_{2}(s_{1})&T_{2}(s_{2})&T_{2}(s_{3})&\dots&\frac{1}{2}T_{2}(s_{m_{\ell}})\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \frac{1}{4}T_{m-1}(s_{1})&\frac{1}{2}T_{m-1}(s_{2})&\frac{1}{2}T_{m-1}(s_{3})&\dots&\frac{1}{4}T_{m-1}(s_{m_{\ell}})\end{pmatrix},$ (3.6)
then somewhat surprisingly, by [38, Equations 6.27-6.28], the Chebyshev coefficient tensor ${\bf C}$ can be computed by the DCT of the function evaluation tensor ${\bf T}$ as
${\bf C}={\bf T}\times_{1}W^{(1)}\times_{2}W^{(2)}\times_{3}W^{(3)}.$ (3.7)
[…]
${\bf C}\approx\hat{\bf C}=\bm{\beta}\times_{1}V^{(1)}\times_{2}V^{(2)}\times_{3}V^{(3)},\quad[\bm{\beta},V^{(1)},V^{(2)},V^{(3)}]=\texttt{tuck\_als}({\bf C},\varepsilon)$ (3.8)
with $\bm{\beta}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$ and $V^{(\ell)}\in\mathbb{R}^{m_{\ell}\times r_{\ell}}$, $\ell=1,2,3$. See Remark 3.1 for the discussions on Tucker-ALS and other Tucker decomposition algorithms. Substituting ${\bf C}$ in Eq. 3.1 by its Tucker format Eq. 3.8 gives the ChebTuck approximation of $f$, denoted by $\hat{f}_{\bf m}$. In view of Proposition 2.5, the ChebTuck approximation $\hat{f}$ of $f$ can be written as
$f(x_{1},x_{2},x_{3})\approx\tilde{f}_{\bf m}(x_{1},x_{2},x_{3})\approx\hat{f}_{\bf m}(x_{1},x_{2},x_{3})=\bm{\beta}\times_{1}v^{(1)}(x_{1})\times_{2}v^{(2)}(x_{2})\times_{3}v^{(3)}(x_{3}).$ (3.9)
where $v^{(\ell)}_{j_{\ell}}(x_{\ell})=\sum_{i_{\ell}=1}^{m_{\ell}}V^{(\ell)}_{i_{\ell},j_{\ell}}T_{i_{\ell}-1}(x_{\ell})$ for $j_{\ell}=1,\cdots,r_{\ell}$, $\ell=1,2,3$.
@@
1:Given a function $f:[-1,1]^{3}\to\mathbb{R}$, the Chebyshev degrees ${\bf m}=(m_{1},m_{2},m_{3})$ and tuck_als tolerance threshold $\varepsilon>0$.
@@
3:Compute the function evaluation tensor ${\bf T}$ using Eq. 3.5.
4:Compute the Chebyshev coefficient tensor ${\bf C}$ using Eq. 3.7.
5:Compute the Tucker decomposition: $[\bm{\beta},V^{(1)},V^{(2)},V^{(3)}]=\texttt{tuck\_als}({\bf C},\varepsilon)$.
6:Form the factor functions $v^{(\ell)}_{j_{\ell}}(x_{\ell})=\sum_{i_{\ell}=1}^{m_{\ell}}V^{(\ell)}_{i_{\ell},j_{\ell}}T_{i_{\ell}-1}(x_{\ell})$ for $\ell=1,2,3$.
7:Return $\hat{f}_{\bf m}(x_{1},x_{2},x_{3})=\bm{\beta}\times_{1}v^{(1)}(x_{1})\times_{2}v^{(2)}(x_{2})\times_{3}v^{(3)}(x_{3})$.
@@
It is important to note that tuck_als in 4 of Algorithm 1 can be any Tucker decomposition algorithm that can approximate a tensor in Tucker format with almost optimal rank parameters. Tucker-ALS [14, 15] is the classical algorithm for this purpose and is widely used in practice. It enjoys strong theoretical guarantees such that for any given $\varepsilon>0$, it can approximate the original tensor in the sense of relative error $\|{\bf C}-\hat{\bf C}\|_{\mathrm{F}}\leq\varepsilon\|{\bf C}\|_{\mathrm{F}}$ with almost optimal rank parameters, where $\|\cdot\|_{\mathrm{F}}$ is the Frobenius norm of a tensor. However, it is computational expensive since it requires the access to all entries of the tensor ${\bf C}$, amounting to $O(m^{3})$ space complexity and successive SVDs of the unfolding matrices, amounting to $O(m^{4})$ algebraic operations. Alternatives to the classical Tucker-ALS are, e.g., the multigrid algorithm [31], cross-based algorithms [40, 16, 42, 22], or randomized algorithms [11, 32]. In case of the relatively small rank of the input tensor, those algorithms are more efficient in terms of computational cost and memory usage, but they may contain several heuristics, especially the cross-based algorithms with no theoretical accuracy guarantees for given $\varepsilon$. The discussions on the difference and choice of Tucker decomposition algorithms are beyond the scope of this paper. We use the classical Tucker-ALS algorithm in our theoretical analysis and numerical experiments demonstrated reliable efficiency for low-rank approximation of the coefficients tensor ${\bf C}$.

### v2

@@ 2.2 Transforming hybrid representation to the standard Tucker format
Let the two-level orthogonal Tucker tensor ${\bf A}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ be given in the form Eq. 6, where the core $\bm{\beta}$ is parametrized by the rank ${\bf r}^{0}=(r_{1}^{0},r_{2}^{0},r_{3}^{0})$ orthogonal Tucker tensor as in Eq. 5. Then $\bf A$} takes the form of a rank ${\bf r}^{0}$ orthogonal Tucker representation with the core tensor $\bm{\eta}\in\mathbb{R}^{r_{1}^{0}\times r_{2}^{0}\times r_{3}^{0}}$ and the orthogonal Tucker factor matrices $A^{(\ell)}G^{(\ell)}\in\mathbb{R}^{n_{\ell}\times r_{\ell}^{0}}$, i.e.,
(9) ${\bf A}=\bm{\eta}\times_{1}(A^{(1)}G^{(1)})\times_{2}(A^{(2)}G^{(2)})\times_{3}(A^{(3)}G^{(3)}).$
Proof 2.5.
Substituting $\bm{\beta}$ Eq. 5 into Eq. 6 and applying contracted product operations to the resultant Tucker tensor proves Eq. 9. The orthogonality of the factor matrices follows from $(G^{(\ell)})^{\top}(A^{(\ell)})^{\top}A^{(\ell)}G^{(\ell)}=I_{r_{\ell}^{0}}$.
Notice that the resultant conventional Tucker tensor (9) can be easily transformed to the canonical tensor format by using the so-called Tuck-to-can transform applied to the core tensor $\bm{\eta}$, see [39, Def. 2.3, (2.6)], [38, Rem. 2.7], and [34].
The following statement is important in further use of the hybrid functional representation in multilinear algebra, see also (17).
@@
The format Eq. 7 can be converted to the standard functional Tucker format as in Definition 2.1, i.e., Eq. 7 can be equivalently written as
(10) $f(x_{1},x_{2},x_{3})=\bm{\beta}\times_{1}v^{(1)}(x_{1})\times_{2}v^{(2)}(x_{2})\times_{3}v^{(3)}(x_{3}),$
where the Tucker factor functions are given by $v^{(\ell)}(x_{\ell})=T_{1:m_{\ell}}(x_{\ell})V^{(\ell)}\in\mathbb{R}^{r_{\ell}}$, i.e., each entry of $v^{(\ell)}$ is a linear combination of the Chebyshev polynomials, while the core tensor is given by Eq. 8. Furthermore, if Eq. 8 is an orthogonal Tucker tensor, then $\{T_{1:m_{\ell}}(x_{\ell})V^{(\ell)}\}$ is a set of $r_{\ell}^{0}$ orthogonal polynomials with the inner product induced by the orthogonality property of Chebyshev polynomials.
The proof is in essence the same as Proposition 2.4, i.e., follows from the easy calculation of contracted products. We nevertheless write out several useful equivalent representations of the ChebTuck format explicitly for further use.
Proof 2.7.
From Eq. 7 and Eq. 8, we have
(11) $f(x_{1},x_{2},x_{3})=\sum_{i_{1}=1}^{m_{1}}\sum_{i_{2}=1}^{m_{2}}\sum_{i_{3}=1}^{m_{3}}{\bf C}_{i_{1},i_{2},i_{3}}T_{i_{1}}(x_{1})T_{i_{2}}(x_{2})T_{i_{3}}(x_{3})$
$=$ $\sum_{j_{1}=1}^{r_{1}}\sum_{j_{2}=1}^{r_{2}}\sum_{j_{3}=1}^{r_{3}}\beta_{j_{1},j_{2},j_{3}}\sum_{i_{1}=1}^{m_{1}}V^{(1)}_{i_{1},j_{1}}T_{i_{1}}(x_{1})\sum_{i_{2}=1}^{m_{2}}V^{(2)}_{i_{2},j_{2}}T_{i_{2}}(x_{2})\sum_{i_{3}=1}^{m_{3}}V^{(3)}_{i_{3},j_{3}}T_{i_{3}}(x_{3})$
@@
where $v^{(\ell)}_{j_{\ell}}(x_{\ell})=\sum_{i_{\ell}=1}^{m_{\ell}}V^{(\ell)}_{i_{\ell},j_{\ell}}T_{i_{\ell}}(x_{\ell})$ for $\ell=1,2,3$. The desired result then follows. The orthogonality property is based on the argument analogues to that of the algebraic case but reformulated for polynomial arrays. We note that the inner product should be properly reweighted.
@@
We can see that the ChebTuck format is a degree ${\bf m}=(m_{1},m_{2},m_{3})$ polynomial in Chebyshev basis, where the coefficients are given as a Tucker tensor. For this reason, we sometimes also write a subscript $\bf m$} to denote the ChebTuck format of certain degrees, e.g., $f_{\bf m}$.
3 Numerical methods for hybrid Chebyshev-Tucker approximation of trivariate functions
In what follows, we introduce the numerical methods for rank-structured approximation of trivariate functions in the novel ChebTuck tensor format introduced in Definition 2.3. The basic idea is to first compute the Chebyshev coefficient tensor $\bf C$}, either in full tensor format or in some low-rank format, and then convert $\bf C$} to the desired Tucker format with quasi-optimal $\varepsilon$-rank. Our techniques can be applied directly to multivariate functions with moderate regularity for which the Chebyshev tensor product interpolation demonstrates satisfactory convergence rates. Furthermore, we extend the presented approach beyond the class of regular functions and successfully apply it to functions with multiple singularities by using the range-separated tensor decomposition.
@@ 3.1.1 Functional case
Given $f:[-1,1]^{3}\to\mathbb{R}$, we consider its approximation by the tensor product of Chebyshev polynomials in the form (c.f. Eq. 7)
(12) $f(x_{1},x_{2},x_{3})\approx\tilde{f}_{\bf m}(x_{1},x_{2},x_{3})=\sum_{i_{1}=1}^{m_{1}}\sum_{i_{2}=1}^{m_{2}}\sum_{i_{3}=1}^{m_{3}}{\bf C}_{i_{1},i_{2},i_{3}}T_{i_{1}}(x_{1})T_{i_{2}}(x_{2})T_{i_{3}}(x_{3}),$
where $T_{k}(x)$ is the (k1)(k-1)-st Chebyshev polynomial, ${\bf C}\in\mathbb{R}^{m_{1}\times m_{2}\times m_{3}}$ is the Chebyshev coefficient tensor (CCT).
There are mainly two ways to view the approximant $\tilde{f}_{\bf m}$ which lead to two different methods to compute the CCT $\bf C$}. The first way is to view $\tilde{f}_{\bf m}$ as a truncation of the tensor product Chebyshev series expansion of $f$ and compute the coefficients by projection and integration. In this case, the Chebyshev coefficients are in fact straightforward to derive and given by the $w$-weighted inner product of $f$ and the corresponding Chebyshev basis, i.e., ${\bf C}_{i_{1},i_{2},i_{3}}=\int_{[-1,1]^{3}}f(\mathbf{x})T_{i_{1}}(x_{1})T_{i_{2}}(x_{2})T_{i_{3}}(x_{3})w(\mathbf{x})\operatorname{d}\mathbf{x}$, where $w(\mathbf{x})=\prod_{\ell=1}^{3}(1-x_{\ell}^{2})^{-1/2}$ is the Chebyshev weight function, see e.g., [45, 46], but the computation is troublesome since one needs to evaluate those 3D integrals. Here, we used the notation $\mathbf{x}=(x_{1},x_{2},x_{3})$. The second way is to view $\tilde{f}_{\bf m}$ as an interpolation of $f$ at Chebyshev nodes, namely, $\tilde{f}_{\bf m}$ satisfies the interpolatory property $\tilde{f}_{\bf m}(s_{i_{1}}^{(1)},s_{i_{2}}^{(2)},s_{i_{3}}^{(3)})=f(s_{i_{1}}^{(1)},s_{i_{2}}^{(2)},s_{i_{3}}^{(3)})$ for all $i_{\ell}=1,\ldots,m_{\ell}$, $\ell=1,2,3$, where $s_{i_{\ell}}^{(\ell)}$ are the Chebyshev nodes of the second kind on each mode given by $s_{i_{\ell}}^{(\ell)}=\cos\left((i_{\ell}-1)\pi/(m_{\ell}-1)\right)$. The derivation of the CCT $\bf C$} in this case is more involved while it turns out that the computation can be done efficiently by the discrete cosine transform (DCT). The accuracy of these two approaches is in practice similar [45] and we will therefore focus on the interpolation approach in this paper.
@@
(13) ${\bf T}_{i_{1},i_{2},i_{3}}=f(s_{i_{1}}^{(1)},s_{i_{2}}^{(2)},s_{i_{3}}^{(3)}),\quad i_{\ell}=1,\ldots,m_{\ell},\quad\ell=1,2,3$
and the inverse DCT matrix $W^{(\ell)}\in\mathbb{R}^{m_{\ell}\times m_{\ell}}$ as
(14) $W^{(\ell)}=\dfrac{2}{m_{\ell}-1}\begin{pmatrix}\frac{1}{4}T_{1}(s_{1})&\frac{1}{2}T_{1}(s_{2})&\frac{1}{2}T_{1}(s_{3})&\dots&\frac{1}{4}T_{1}(s_{m_{\ell}})\\ \frac{1}{2}T_{2}(s_{1})&T_{2}(s_{2})&T_{2}(s_{3})&\dots&\frac{1}{2}T_{2}(s_{m_{\ell}})\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \frac{1}{4}T_{m_{\ell}}(s_{1})&\frac{1}{2}T_{m_{\ell}}(s_{2})&\frac{1}{2}T_{m_{\ell}}(s_{3})&\dots&\frac{1}{4}T_{m_{\ell}}(s_{m_{\ell}})\end{pmatrix}$,
then by [47, Equations 6.27—6.28], the CCT $\bf C$} can be computed by the DCT of the function evaluation tensor $\bf T$} as
(15) ${\bf C}={\bf T}\times_{1}W^{(1)}\times_{2}W^{(2)}\times_{3}W^{(3)}.$
Given an error tolerance $\varepsilon>0$, we can approximate the coefficient tensor $\bf C$} by a Tucker tensor with almost optimal rank parameter ${\bf r}=(r_{1},r_{2},r_{3})$ using the Tucker-ALS algorithm denoted by $\texttt{tuck\_als}({\bf C},\varepsilon)$
(16) ${\bf C}\approx\hat{\bf C}=\bm{\beta}\times_{1}V^{(1)}\times_{2}V^{(2)}\times_{3}V^{(3)},\quad[\bm{\beta},V^{(1)},V^{(2)},V^{(3)}]=\texttt{tuck\_als}({\bf C},\varepsilon)$
with $\bm{\beta}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$ and $V^{(\ell)}\in\mathbb{R}^{m_{\ell}\times r_{\ell}}$, $\ell=1,2,3$. See Remark 3.1 for the discussions on Tucker-ALS and other Tucker decomposition algorithms. Substituting $\bf C$} in Eq. 12 by its Tucker format Eq. 16 gives the ChebTuck approximation of $f$, denoted by $\hat{f}_{\bf m}$. In view of Proposition 2.6, the ChebTuck approximation $\hat{f}$ of $f$ can be written as
(17) $f(\textbf{x})\approx\tilde{f}_{\bf m}(x_{1},x_{2},x_{3})\approx\hat{f}_{\bf m}(x_{1},x_{2},x_{3})=\bm{\beta}\times_{1}v^{(1)}(x_{1})\times_{2}v^{(2)}(x_{2})\times_{3}v^{(3)}(x_{3}).$
where $v^{(\ell)}_{j_{\ell}}(x_{\ell})=\sum_{i_{\ell}=1}^{m_{\ell}}V^{(\ell)}_{i_{\ell},j_{\ell}}T_{i_{\ell}}(x_{\ell})$ for $j_{\ell}=1,\ldots,r_{\ell}$, $\ell=1,2,3$.
@@
1:Given a function $f:[-1,1]^{3}\to\mathbb{R}$, the Chebyshev degrees ${\bf m}=(m_{1},m_{2},m_{3})$ and tuck_als tolerance $\varepsilon>0$.
@@
3:Compute the function evaluation tensor $\bf T$} by Eq. 13, then the CCT $\bf C$} by Eq. 15.
4:Compute the Tucker decomposition: $\bm{\beta},V^{(1)},V^{(2)},V^{(3)}]=\texttt{tuck\_als}({\bf C},\varepsilon)$ .
5:Form the factor functions $v^{(\ell)}_{j_{\ell}}(x_{\ell})=\sum_{i_{\ell}=1}^{m_{\ell}}V^{(\ell)}_{i_{\ell},j_{\ell}}T_{i_{\ell}}(x_{\ell})$ for $\ell=1,2,3$.
6:Return $\hat{f}_{\bf m}(x_{1},x_{2},x_{3})=\bm{\beta}\times_{1}v^{(1)}(x_{1})\times_{2}v^{(2)}(x_{2})\times_{3}v^{(3)}(x_{3})$.
@@
Notice that tuck_als in line 4 of Algorithm 1 implementing the classical Tucker-ALS algorithm [18, 19] for this purpose, is widely used in practice for moderate $d$ and $m$. It enjoys strong theoretical guarantees such that for any given $\varepsilon>0$, it approximates the original tensor in the sense of relative error $\|{\bf C}-\hat{\bf C}\|_{\mathrm{F}}\leq\varepsilon\|{\bf C}\|_{\mathrm{F}}$ with almost optimal rank parameters, where $\|\cdot\|_{\mathrm{F}}$ is the Frobenius norm of a tensor. However, it is computational expensive since it requires the access to all entries of the tensor $\bf C$}, amounting to $O(m^{3})$ space complexity and successive SVDs of the unfolding matrices, amounting to $O(m^{4})$ algebraic operations. Alternatives to the classical Tucker-ALS are, e.g., the multigrid algorithm [39], cross-based algorithms [49, 21, 20, 50, 29], or randomized algorithms [15, 40]. In case of the relatively small rank of the input tensor, those algorithms are more efficient in terms of computational cost and memory usage. In particular, using the cross-based algorithms can avoid the evaluation of the full tensor $\bf T$}. However, they contain several heuristics, especially the cross-based algorithms have no theoretical accuracy guarantees for given $\varepsilon$. The discussions on the difference and choice of Tucker decomposition algorithms are beyond the scope of this paper. We use the classical Tucker-ALS algorithm in our theoretical analysis and numerical experiments demonstrated reliable efficiency for low-rank approximation of the coefficient tensor $\bf C$}.